OpenStudy (anonymous):

help me with this I am stuck please help me finish I just need an explanation please. the population of a city is growing at a rate proportional to its population. the population 20 years ago was 100 000 and today it is 150 000. find the population 20 years from now

OpenStudy (anonymous):

I need to apply logarithms application to this problem.

OpenStudy (anonymous):

From a calculus perspective, you can start with \[dP = \alpha Pdt\] where P(t) is the population This will give you something like \[P(t) = Ce^{\alpha t}\] with C and a being constants. Consider using "20 years ago" as t=T. Then "today" is t=T+20. Plugging these in you have two equations:\[100,000=Ce^{\alpha T}\]\[150,000 = Ce^{\alpha(T+20)}\] If you divide the second equation by the first, you get: \[1.5 = e^{\alpha 20}\] and you can find a. Then, do something similar with "twenty years from now" (t=T+40!)... \[100,000 = Ce^{\alpha T}\]\[X = Ce^{\alpha (T+40)}\] Divide the second by the first: \[\frac{X}{100,000} = e^{\alpha 40}\] Where you know a! Solve for your answer X.

OpenStudy (anonymous):

can you help please

OpenStudy (anonymous):

Which part doesn't make sense?

OpenStudy (anonymous):

150 000=ce^at

OpenStudy (anonymous):

I have never seen that can you use the general form of A=Poe^kt

OpenStudy (anonymous):

You have seen \[A = P_0e^{kt}\] right? So we re-word it a bit. :) But let's call "twenty years ago" t = 0, okay?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then you can solve for P_0

OpenStudy (anonymous):

So what is P_0?

OpenStudy (anonymous):

Po=100 000

OpenStudy (anonymous):

Okay, so now we have \[A = (100,000)e^{kt}\] If twenty years ago was t=0, what does t equal today?

OpenStudy (anonymous):

is it A_o or A

OpenStudy (anonymous):

One step at at a time. t is the time in years, right? So, t=? today?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

Okay, so now we have (for today) \[A = (100,000)e^{20a}\] AND we know that the population today is 150,000. Where does population go into that equation?

OpenStudy (anonymous):

goe to A

OpenStudy (anonymous):

Great. So you have \[150,000 = 100,000e^{20a}\] And you can solve for a. So now, find a.

OpenStudy (anonymous):

ok doing it now just a sec

OpenStudy (anonymous):

got 0.0202

OpenStudy (anonymous):

Okay. Good. Now you know a (or k) and P_0. So your population equation is:\[A = (100,000)e^{0.0202t}\] Finally, twenty years from now, t=?

OpenStudy (anonymous):

will it be 20 again?

OpenStudy (anonymous):

Not quite. t=0 was twenty years ago. t=20 is today. t=? is in another twenty years, so t=?

OpenStudy (anonymous):

40

OpenStudy (anonymous):

Yep. So plug in t=40, and find A.

OpenStudy (anonymous):

I got 100 000e^0.808

OpenStudy (anonymous):

Good. Keep going

OpenStudy (anonymous):

(Plug in e so you get the final answer)

OpenStudy (anonymous):

I did and the final result is 14.2486...

OpenStudy (anonymous):

Hmm, that sounds a little low. What is e^0.8?

OpenStudy (anonymous):

yes did to me too

OpenStudy (anonymous):

I have been making mistakes when I get to this step

OpenStudy (anonymous):

Are you using a calculator on the computer or a handheld calculator?

OpenStudy (anonymous):

handheld

OpenStudy (anonymous):

Which model?

OpenStudy (anonymous):

(Once you get the gist of it, you can practice, and check your answers using Google. If you type e^0.8 into Google, you will get the right answer!)

OpenStudy (anonymous):

casio scientific calculator

OpenStudy (anonymous):

i dont know if i need to divide or take log of natural number

OpenStudy (anonymous):

For e^0.8, you are raising the number e (~2.7...) to the 0.8 power. If you calculator looks like this: http://www.ivgstores.com/prodimages-cdls/pet-casi/CIOFX115MS-L.jpg then try pressing "shift" then "ln" then 0.8.

OpenStudy (anonymous):

ln 0.8/ln 100 000 ?

OpenStudy (anonymous):

ln is the opposite of e. Think of it like ln takes the square root of a number, and e squares a number.

OpenStudy (anonymous):

so sorry I dont understand what you mean

OpenStudy (anonymous):

Let's just say that you have a number 2. You can multiply 2 by 4. You can divide 2 by 4. Those do the opposite things, right? Well, logs (ln) and exponentials (e) also do the opposite things.

OpenStudy (anonymous):

In this way: \[ln(2) = 0.69\] and \[e^{0.69} = 2\]

OpenStudy (anonymous):

And from this, you get properties like \[\ln{(e^x)} = x\] and \[e^{\ln(x)} = x\] (Which is like saying "If I do multiplication by 4 and division by 4 to the number 2, I get the number 2 back!)

OpenStudy (anonymous):

I still get a lower number

OpenStudy (anonymous):

Try it in Google first. What does google say is e^0.8?

OpenStudy (anonymous):

=2.24341666

OpenStudy (anonymous):

Okay, so then \[ 100,000e^{0.8} = 224,342\] Which is reasonable. Now you just need to figure out how to get 2.2 in your calculator for e^0.8

OpenStudy (anonymous):

I did shift e^0.808

OpenStudy (anonymous):

Did the display in your calculator say e^0.808? If so, then you should get 2.24 in your caculator

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good. Now do you have your answer?

OpenStudy (anonymous):

wow what a day. i just have to do 100 000e^0.08

OpenStudy (anonymous):

Well, not e^0.08! Make sure you do 100 000e^0.808

OpenStudy (anonymous):

oh yeah I got 224.3416664

OpenStudy (anonymous):

224,312 people sounds reasonable if today there are 150,000 and twenty years ago there were 100,000 people.

OpenStudy (anonymous):

yes. your great thanks for your patience and your help

OpenStudy (anonymous):

no problem. keep working with your calculator. keep practicing. it will get more natural!

OpenStudy (anonymous):

I will